If you know the area in square feet of a rectangle but don't know what the perimeter is how can you figure that out? I know that you can use the formula 2L + 2W, but I am trying to figure out how 300 ft of fencing can go around a rectangle that has an area of 5000 square ft.
Answer by arunpaul(104)

let the side of rectangle are L and B
as per question L*B = 5000 sft
again 2(L+B) = 300 - eq1
again L = 5000/B
then put the value in the equation 1
then 2((5000/B)+B)= 300
or (10000/B)+2B = 300
or (10000/B) = 300-2B
or 10000= 300B -2B^2
or 5000=150B -B^2
solve the equation by trial and error or by the equation
x1,2 = -b ± sqrt( b2-4ac )
------------------
2a
here the equation is
B^2-150B+5000 = 0
so a= 1
b= 150
c=5000
putting thevalue
u can get the solution for B
such as
((-150)+sqrt(150^2-4*1*5000))/(2*1)
((-150)-sqrt(150^2-4*1*5000))/(2*1)
solving the above equation we get B = -10000 and -10150
so no real solution
that means your question is wrong
check my answer and mail again